# Fishers Harvest Parallel Unlearning in Inherited Model Networks

### v1

@@ 5.2. Learning Framework Setup and Benchmarks
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
@@
AlexNet 1 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 98.17 95.98 91.32 91.40 86.24 83.99 20.31 40.49 42.36 0.09 0.11 0.11
${AD}_{f}\downarrow$ 96.66 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 96.88 96.68 93.16 82.36 80.47 79.99 21.20 40.56 41.92 0.08 0.13 0.11
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 94.67 96.45 93.45 71.59 71.43 76.06 20.30 39.76 40.60 0.09 0.11 0.11
${AD}_{f}\downarrow$ 96.61 98.30 98.30 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 97.16 95.61 92.20 99.99 99.99 99.99 75.97 69.56 76.15 19.42 39.46 39.35 0.08 0.13 0.11
${AD}_{f}\downarrow$ 98.57 99.28 99.28 0.00 0.00 0.00 20.28 20.28 20.28
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.09 93.51 95.22 22.21 41.52 42.93 0.15 0.24 0.24
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 85.89 94.89 95.79 21.19 41.72 43.20 0.17 0.23 0.26
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 80.20 94.82 97.03 22.03 42.26 43.25 0.17 0.25 0.25
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 75.36 98.07 98.96 19.10 42.60 41.36 0.16 0.27 0.27
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 5. Distributed Data-Parallel Unlearning Performance On CIFAR-100
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
@@
AlexNet 1 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 97.23 98.85 97.66 30.13 31.26 31.14 0.09 0.14 0.23
${AD}_{f}\downarrow$ 99.99 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.96 99.96 99.96 99.96 99.96 99.96 96.14 97.07 96.71 26.43 26.47 26.41 0.11 0.16 0.27
${AD}_{f}\downarrow$ 99.96 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 94.58 90.41 90.97 24.33 24.36 24.37 0.10 0.14 0.23
${AD}_{f}\downarrow$ 99.09 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 91.53 76.19 80.70 21.03 22.96 22.10 0.10 0.14 0.24
${AD}_{f}\downarrow$ 97.15 99.96 99.96 0.00 0.00 0.00 9.73 0.00 11.08
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.87 94.67 93.67 32.53 31.95 32.18 0.30 0.34 0.64
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
2 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.94 94.58 92.40 28.28 27.22 28.47 0.30 0.34 0.66
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 86.44 93.95 89.26 91.77 25.37 24.97 25.13 0.31 0.35 0.66
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 99.96 90.44 95.79 94.19 22.60 21.92 24.19 0.34 0.34 0.68
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 9.73 0.00 0.00
Table 6. Federated Unlearning Performance On TinyImageNet With Multiple Label Distribution
Model #flimit-$\#\cap_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 96.39 84.75 77.89 3000.66 1998.46 994.32 4.69 2.39 2.30
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.50 99.99 99.80 99.51 99.99 99.99 96.64 85.40 85.00 3009.53 1010.45 2011.14 5.12 2.69 2.14
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 96.22 81.49 78.83 3022.64 2020.67 1021.34 4.98 2.64 2.59
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.71 96.01 83.04 79.15 3039.64 2041.32 1046.14 4.96 2.14 2.50
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.23 98.54 98.44 349.32 246.43 123.74 3.16 1.65 1.68
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.98 99.44 99.49 355.32 247.64 128.43 2.97 1.46 1.59
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 94.43 98.86 99.38 354.83 246.43 129.35 2.89 1.54 1.38
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.36 99.34 99.32 351.38 244.34 125.32 3.10 1.69 1.39
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 7. Distributed Data-Parallel Unlearning Performance On CIFAR-100 With Multiple Label Distribution
Model #flimit-$\#\cap_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161 60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 95.18 97.07 99.99 999.41 996.43 997.64 0.15 0.08 0.09
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.80 99.51 99.99 99.99 90.53 97.07 98.82 1004.64 1006.31 1006.14 0.19 0.09 0.11
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 95.29 97.07 99.89 1019.43 1020.16 1018.64 0.18 0.07 0.10
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 87.15 97.07 96.36 1031.43 1029.64 1030.64 0.18 0.10 0.11
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.62 88.94 87.87 30.36 33.18 37.31 0.60 0.32 0.29
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.50 89.35 87.89 28.30 29.13 52.33 0.59 0.34 0.30
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 95.31 88.17 93.06 26.39 23.37 56.61 0.58 0.28 0.30
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
0% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 88.87 88.86 89.16 22.03 24.36 46.42 0.64 0.37 0.32
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
5.3. Label Category Settings
The unlearned label categories are represented as #$\#C_{f}$. The distribution of labels among different models may also vary in different data distributions. For instance, model $w_{a}$ aims to unlearn labels 1, 2, 3, and 4, while model $w_{b}$ intends to unlearn labels 2, 3, 5, and 6. In this case, there is an overlap between the unlearned label categories of model $w_{a}$ and model $w_{b}$, specifically labels 2 and 3. We refer to the degree of overlapping among these unlearned label categories as #flimit-$\#\cap_{f}$.
The unlearning process is applied to individual categories and various combinations of labels. For the CIFAR-100 dataset, the combinations we select include 1, 2, 4, and 10 categories. For the TinyImageNet dataset, the selected combinations include 1, 4, 6, and 10 categories. The overlap of the unlearned label categories is divided into different proportions, including 60%, 40%, 20%, and 0%.
6. Evaluation
In this section, we use NVIDIA 4090 GPU to set up a test platform that demonstrates the impact of the FIUn method on the accuracy of unlearned and retained labels in four frameworks: FL, DDPL, IL and TL. When assessing accuracy and time consumption across these learning frameworks during the model unlearning process, we compare our proposed FIUn method with existing unlearning methods and use re-training as a benchmark. We also analyze the impact of the inheritance depth and parameter layers of the models on unlearning effectiveness.
@@ 6. Evaluation
Figure 7. Overlap #$\cap_{f}$ analysis, #$C_{f}$=15.
6.1. General Label Unlearning Analysis
We consider existing emerging unlearning methods for comparison: 1) Fine-tuning, which adjusts the original model using the remaining dataset (Odusami et al., 2021), 2) Gradient ascent (GA), employing negative gradients for unlearning (Zhang et al., 2024), and 3) Distill, which distills knowledge from the original model into a student model using the remaining data (Sinha et al., 2023).
Performance comparisons are depicted in Figure 5 for unlearning accuracy $\bigtriangleup_{acc}$ and detailed in Table 1 for unlearning speed. The results demonstrate that our method not only enhances unlearning speed by up to 99% over alternative unlearning methods but also consistently outperforms them in unlearning accuracy.
These results indicate that our method exhibits outstanding effectiveness and efficiency in handling the complexity of model unlearning. In what follows, we evaluate the proposed parallel unlearning method in comparison with the most robust method—re-training. Within all considered learning frameworks, our proposed FIUn method achieves a maximum increase in unlearning speed of 99% compared to the re-training benchmark across different datasets and models. The detailed results are summarized in Tables 2 to 5, with more experiments provided in Appendix .2.1.
@@ 6.1. General Label Unlearning Analysis
1) Single-label unlearning. For all frameworks and model types in CIFAR100, experimental results show that the $AD_{f}$ metric reaches 0, strongly proving that our method effectively unlearns single-class labels. Meanwhile, the average $AD_{r}$ metric is as high as 94.53%, further confirming our method’s efficiency in retaining labels. For all frameworks and model types in TinyImageNet, the $AD_{f}$ metric also reaches 0, and the average $AD_{r}$ metric reaches 79.49%. This significantly validates the effectiveness of our method.
@@
2) Multi-label unlearning. For the CIFAR100 dataset, with #$C_{f}=2,4,10$, the average $AD_{f}$ metric across all frameworks and model types is 1.93%, while the average $AD_{r}$ metric reaches 86.01%. For the TinyImageNet dataset, the average $AD_{f}$ metric is 0.21%, and the average $AD_{r}$ metric is 83.54%. These results further demonstrate the effectiveness of our method. On the other hand, for both the IL and TL frameworks, as the number of unlearning label classes increases, the $AD_{f}$ metric for all models approaches 0. However, the $AD_{r}$ performance of the AlexNet and DenseNet161 models shows a downward trend, while the $AD_{r}$ performance of the ResNet18 model remains stable.
@@
Further observation reveals that the number of parameters in the last layer of DenseNet161 is 4.3 times that of ResNet18, and the number of parameters in the last layer of AlexNet is 7.9 times that of ResNet18. These results indicate that as the number of model parameters and the classes of forgotten labels increases, the $AD_{r}$ performance may decline. This further suggests that while our proposed method effectively unlearns specified labels, it can somewhat affect the accuracy of retained labels.
(a) FL TinyImageNet by ResNet18
(b) FL CIFAR100 by ResNet18
(c) FL TinyImageNet by DenseNet161
(d) FL CIFAR100 by DenseNet161
(e) IL TinyImageNet by ResNet18
(f) IL CIFAR100 by ResNet18
(g) IL TinyImageNet by DenseNet161
(h) IL CIFAR100 by DenseNet161
Figure 8. Inheritance depth analysis #$C_{f}=4$.
(a) FL AlexNet in CIFAR100
(b) FL CIFAR100 by ResNet18
(c) IL CIFAR100 by ResNet18
(d) IL CIFAR100 by AlexNet
(e) FL TinyImageNet by DenseNet161
(f) FL TinyImageNet by ResNet18
(g) IL TinyImageNet by ResNet
(h) IL TinyImageNet by DenseNet161

### v4

@@ VII-A General Label Unlearning Analysis
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
@@
DenseNet161
1 $AD_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 88.03 77.64 76.02 999.41 1996.43 2997.64
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 2.51 4.59 4.95
4 $AD_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.01 71.43 71 1004.64 2006.31 3006.14
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 3.16 4.69 4.89
6 $AD_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 80.00 67.67 66.58 1019.43 2020.16 3018.64
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 3.10 4.96 4.75
ResNet18
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.73 99.97 99.46 112.95 216.39 314.29
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.35 3.68 3.42
4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.17 99.98 99.98 109.89 209.31 310.42
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.63 3.26 3.57
6 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 99.54 99.98 99.98 107.39 208.12 305.32
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.46 3.37 3.51
TABLE VI: Incremental Unlearning Performance on TinyImageNet
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
@@
DenseNet161
1 $AD_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 89.76 78.50 73.87 991.43 1995.36 3989.47
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 2.75 4.85 4.43
4 $AD_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.98 65.87 59.11 1006.32 2004.75 4009.41
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 2.64 4.43 4.64
6 $AD_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 86.81 61.02 55.04 1019.43 2019.50 4019.46
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 2.36 4.47 4.74
ResNet18
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.66 96.33 96.07 120.32 221.75 439.85
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.37 3.75 3.43
4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.84 96.34 96.70 115.32 217.33 435.17
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.41 3.34 3.64
6 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 98.82 99.99 98.59 110.32 211.35 431.37
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 1.47 3.46 3.75
@@
Figure 8: Overlap $\cap_{f}$ analysis, $C_{f}$=15. MFIM stands for Merging FIM function.
@@
1) Single-label unlearning. For all frameworks and model types in CIFAR100, experimental results show that the $AD_{f}$ metric reaches 0, strongly proving that our method effectively unlearns single-class labels. Meanwhile, the average $AD_{r}$ metric is as high as 94.53%, further confirming our method’s efficiency in retaining labels. For all frameworks and model types in TinyImageNet, the $AD_{f}$ metric also reaches 0, and the average $AD_{r}$ metric reaches 79.49%. This validates the effectiveness of our method.
@@
2) Multi-label unlearning. For the CIFAR100 dataset, with $C_{f}=2,4$, the average $AD_{f}$ metric across all frameworks and model types is 1.93%, while the average $AD_{r}$ metric reaches 86.01%. For the TinyImageNet dataset, the average $AD_{f}$ metric is 0.21%, and the average $AD_{r}$ metric is 83.54%. These results further demonstrate the effectiveness of our method. On the other hand, for both the IL and TL frameworks, as the number of unlearning label classes increases, the $AD_{f}$ metric for all models approaches 0. However, the $AD_{r}$ performance of the AlexNet and DenseNet161 models shows a downward trend, while the $AD_{r}$ performance of the ResNet18 model remains stable.
TABLE VII: Distributed Data-Parallel Unlearning Performance on CIFAR100 with Multiple Label Distribution
Model $\cap_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161
60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 95.18 97.07 99.99 999.41 996.43 997.64
$AD_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 0.15 0.08 0.09
40% $AD_{r}\uparrow$ 99.99 99.99 99.80 99.51 99.99 99.99 90.53 97.07 98.82 1004.64 1006.31 1006.14
$AD_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 0.19 0.09 0.11
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 95.29 97.07 99.89 1019.43 1020.16 1018.64
$AD_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 0.18 0.07 0.10
ResNet18
60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.62 88.94 87.87 30.36 33.18 37.31
$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.60 0.32 0.29
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.50 89.35 87.89 28.30 29.13 52.33
$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.59 0.34 0.30
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 95.31 88.17 93.06 26.39 23.37 56.61
$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.58 0.28 0.30
TABLE VIII: Federated Unlearning Performance on TinyImageNet with Multiple Label Distribution
Model $\cap_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161
60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 96.39 84.75 77.89 3000.66 1998.46 994.32
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 4.69 2.39 2.30
40% $AD_{r}\uparrow$ 99.50 99.99 99.80 99.51 99.99 99.99 96.64 85.40 85.00 3009.53 1010.45 2011.14
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 5.12 2.69 2.14
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 96.22 81.49 78.83 3022.64 2020.67 1021.34
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 4.98 2.64 2.59
ResNet18
60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.23 98.54 98.44 349.32 246.43 123.74
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.16 1.65 1.68
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.98 99.44 99.49 355.32 247.64 128.43
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 2.97 1.46 1.59
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 94.43 98.86 99.38 354.83 246.43 129.35
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 2.89 1.54 1.38
@@
Further observation reveals that the number of parameters in the last layer of DenseNet161 is 4.3 times that of ResNet18, and the number of parameters in the last layer of AlexNet is 7.9 times that of ResNet18. These results indicate that as the number of model parameters and the classes of unlearned labels increases, the $AD_{r}$ performance may decline.
Takeaway–faster and better. FIUn outperforms others in unlearning speed by 99%, averaging under one second per model, while effectively removing targeted labels and preserving retained accuracy, matching the effectiveness of the re-training method, which is considered the most effective.
(a) FL, TinyImageNet, ResNet18
(b) FL, CIFAR100, ResNet18
(c) FL, TinyImageNet, DenseNet161
(d) FL, CIFAR100, DenseNet161
(e) IL, TinyImageNet, ResNet18
(f) IL, CIFAR100, ResNet18
(g) IL, TinyImageNet, DenseNet161
(h) IL, CIFAR100, DenseNet161
Figure 9: Inheritance depth analysis $C_{f}=1$.
(a) FL, TinyImageNet, ResNet18
(b) FL, CIFAR100, ResNet18
(c) FL, TinyImageNet, DenseNet161
(d) FL, CIFAR100, DenseNet161
(e) IL, TinyImageNet, ResNet18
(f) IL, CIFAR100, ResNet18
(g) IL, TinyImageNet, DenseNet161
(h) IL, CIFAR100, DenseNet161
Figure 10: Inheritance depth analysis $C_{f}=4$.
(a) FL, AlexNet, CIFAR100
(b) FL, CIFAR100, ResNet18
(c) IL, CIFAR100, ResNet18
(d) IL,CIFAR100, AlexNet
(e) FL, TinyImageNet, DenseNet161
(f) FL, TinyImageNet, ResNet18
(g) IL, TinyImageNet, ResNet18
(h) IL, TinyImageNet, DenseNet161